# Advances in the Shannon Capacity of Graphs

### v1

@@ 8 Outlook
    It was proved in [17] that if the Shannon capacity of a graph is attained at some finite power, then the Shannon capacity of its Mycielskian is strictly larger than that of the original graph. In view of the constructions presented in Section 5, which yield graph families whose Shannon capacity is not attained at any finite power, it would be interesting to determine whether this property also holds in such cases. If it does not, the graph constructions from Section 5 could provide potential candidates for a counterexample.
@@
    By combining Theorems 2.5 and 2.6, the equality $\alpha(\mathsf{G}\boxtimes\mathsf{H})=\alpha(\mathsf{G})\,\alpha(\mathsf{H})$ holds for every simple graph $\mathsf{H}$ if $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$. Moreover, by Theorem 2.12 and Lemma 3.2, if $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$, then $\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})$ holds for all $\mathsf{H}$. From these results, a natural question arises: Is it true that the equality $\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})$ holds for all $\mathsf{H}$ if and only if $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$?. This question was already discussed to some extent in [45], where it was shown that
@@
    while raising the question about the possible gap between the left and right hand sides of (171).
@@
Appendix A An original proof of Theorem 2.16
Theorem A.1.
[16] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@
$\vartheta(\mathsf{G})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}.$ (173)
@@
$\vartheta(\mathsf{H})=\sum_{i=1}^{r}(\mathbf{d}^{\mathrm{T}}\mathbf{v}_{i})^{2}.$ (174)
Assume without loss of generality that the dimensions of $\mathbf{u}_{i},\mathbf{c},\mathbf{v}_{j}$ and $\mathbf{d}$ are $m$ (if the dimensions are distinct, the vectors of the lower dimension can be padded by zeros). Next, let $\mathbf{A}$ as an orthogonal matrix of order $m\times m$ such that:
${\mathbf{A}}{\mathbf{d}}={\mathbf{c}}.$ (175)
Such a matrix $\mathbf{A}$ on the left-hand side of (175), satisfying ${\mathbf{A}}^{\mathrm{T}}{\mathbf{A}}=\operatorname{\mathbf{I}}_{n}$, exists (e.g., the householder matrix defined as ${\mathbf{A}}=\operatorname{\mathbf{I}}_{n}-\frac{2(\mathbf{c}-\mathbf{d})(\mathbf{c}-\mathbf{d})^{\mathrm{T}}}{\|\mathbf{c}-\mathbf{d}\|^{2}}$ provided that $\mathbf{c}\neq\mathbf{d}$, and ${\mathbf{A}}=\operatorname{\mathbf{I}}_{n}$ if $\mathbf{c}=\mathbf{d}$). Let $w=(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ be defined as $\mathbf{w}_{i}=\mathbf{A}\mathbf{v}_{i}$ for every $i\in[r]$. Since the pairwise inner products are preserved under an orthogonal (orthonormal) transformation and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ is an orthonormal representation of $\overline{\mathsf{H}}$, so is $(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$. Likewise, we get
$\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{w}_{i})^{2}=\sum_{i=1}^{r}((\mathbf{A}\mathbf{d})^{\mathrm{T}}(\mathbf{A}\mathbf{v}_{i}))^{2}=\sum_{i=1}^{r}(\mathbf{d}^{\mathrm{T}}\mathbf{v}_{i})^{2}=\vartheta(\mathsf{H}).$ (176)
Next, the representation $(\mathbf{x}_{1},\ldots,\mathbf{x}_{n+r})=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n},\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ is an orthonormal representation of $\overline{\mathsf{G}+\mathsf{H}}$ (since there are no additional nonadjacencies in the graph $\overline{\mathsf{G}+\mathsf{H}}$ in comparison to the disjoint union of the pairs of nonadjacent vertices in $\overline{\mathsf{G}}$ and $\overline{\mathsf{H}}$). Hence, by Theorem 2.14, the equality
$\vartheta(\mathsf{G})+\vartheta(\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{w}_{i})^{2}=\sum_{i=1}^{n+r}(\mathbf{c}^{\mathrm{T}}\mathbf{x}_{i})^{2},$ (177)
@@
$\vartheta(\mathsf{G})+\vartheta(\mathsf{H})\leq\vartheta(\mathsf{G}+\mathsf{H}).$ (178)
@@
$\vartheta(\mathsf{G}+\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{v}_{i})^{2}.$ (179)
@@
$\vartheta(\mathsf{G}+\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{v}_{i})^{2}\leq\vartheta(\mathsf{G})+\vartheta(\mathsf{H}).$ (180)
Combining inequalities (178) and (180) gives the equality in (45). ∎
Appendix B The original proof of Theorem 3.4
@@
Appendix C The original proof of Theorem 5.1
Theorem C.1.
@@
$=\sum_{i=0}^{k}\sum_{j=0}^{k}\binom{k}{i}\binom{k}{j}\alpha(\mathsf{G})^{2k-i-j}\alpha(\mathsf{H}^{i+j}).$
@@
$=\sum_{i=0}^{k}\sum_{j=0}^{k}\binom{k}{i}\binom{k}{j}\alpha(\mathsf{G})^{2k-i-j}\alpha(\mathsf{H}^{i})\,\alpha(\mathsf{G}^{j}).$
Subtracting both equalities gives
$\alpha(\mathsf{K}^{2k})-\alpha(\mathsf{K}^{k})^{2}=\sum_{i=0}^{k}\sum_{j=0}^{k}\left(\binom{k}{i}\binom{k}{j}\alpha(\mathsf{G})^{2k-i-j}\left(\alpha(\mathsf{H}^{i+j})-\alpha(\mathsf{H}^{i})\,\alpha(\mathsf{G}^{j})\right)\right).$
@@
$\alpha(\mathsf{H}^{i+j})-\alpha(\mathsf{H}^{i})\,\alpha(\mathsf{G}^{j})>0.$
Otherwise, for all $m\in\mathbb{N}$, $\alpha(\mathsf{H}^{m})=\alpha(\mathsf{H})^{m}$, so $\Theta(\mathsf{H})=\alpha(\mathsf{H})$ in contradiction to our assumption. Hence $\alpha(\mathsf{K}^{2k})>\alpha(\mathsf{K}^{k})^{2}$ for all $k\geq\max\{i_{0},j_{0}\}$.
Suppose by contradiction that $\Theta(\mathsf{K})=\sqrt[k_{0}]{\alpha(\mathsf{K}^{k_{0}})}$ for some $k_{0}\geq\max\{i_{0},j_{0}\}$. Then $\alpha(\mathsf{K}^{2k_{0}})>\alpha(\mathsf{K}^{k_{0}})^{2}$, which gives
$\Theta(\mathsf{K})\geq\sqrt[2k_{0}]{\alpha(\mathsf{K}^{2k_{0}})}>\sqrt[k_{0}]{\alpha(\mathsf{K}^{k_{0}})}=\Theta(\mathsf{K}),$
thus leading to a contradiction. Hence, it is not attained for any $k\geq\max\{i_{0},j_{0}\}$.
Note that if we assume that $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ for some $k<\max\{i_{0},j_{0}\}$, then for every $n\in\mathbb{N}$
$\Theta(\mathsf{G})\geq\sqrt[nk]{\alpha(\mathsf{K}^{nk})}\geq\sqrt[k]{\alpha(\mathsf{K}^{k})}=\Theta(\mathsf{G}),$
and then, using the same argument for $nk\geq\max\{i_{0},j_{0}\}$, we get a contradiction. Hence, the Shannon capacity of $\mathsf{K}$ is not attained for any $k\in\mathbb{N}$. ∎
Appendix D The original proof of Inequality (159)
Theorem D.1.
@@ Appendix D The original proof of Inequality (159)
$\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}.$ (181)
In addition, inequality (181) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive.
@@
Case 1, if there is an index $t$ that has $u_{t}\in\mathcal{A}$ and $v_{t}\in\mathcal{B}$ then obviously the two vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
Case 2, if there isn’t such an index, then there are two indices $1\leq i,j\leq k$, and two indices $r,s$ between $1$ and $2k$ that have $u_{r}=a_{i}$, $u_{s}=b_{i}$, and $v_{r}=a_{j}$, $v_{s}=b_{j}$. Since $\{a_{i},a_{j}\}$ is an edge in $\mathsf{G}$ if and only if $\{b_{i},b_{j}\}$ is not an edge in $\overline{\mathsf{G}}$, then the vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Hence, $\mathcal{S}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Next, the size(cardinality) of $\mathcal{S}$ is $\binom{2k}{k}n^{k}$, because there are $\binom{2k}{k}$ ways to choose the locations of the coordinates from $\mathcal{A}$ and that dictates the location of the coordinates from $\mathcal{B}$ as well, and, for every coordinate there are $n$ options to choose from, which gives $|\mathcal{S}|=\binom{2k}{k}n^{k}$. Finally, we get
$\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})\geq|\mathcal{S}|=\binom{2k}{k}n^{k},$
@@
$\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq\lim_{k\to\infty}{\sqrt[2k]{\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})}}\geq\lim_{k\to\infty}{\sqrt[2k]{\binom{2k}{k}n^{k}}}=2\sqrt{n}.$
In addition, from [11, Theorem 12], inequality (181) indeed holds with equality when $\mathsf{G}$ is a self-complementary and vertex-transitive graph. ∎

### v2

@@ 8 Outlook
    It was proved in [18] that if the Shannon capacity of a graph is attained at some finite power, then the Shannon capacity of its Mycielskian is strictly larger than that of the original graph. In view of the constructions presented in Section 5, which yield graph families whose Shannon capacity is not attained at any finite power, it would be interesting to determine whether this property also holds in such cases. If it does not, then the graph constructions from Section 5 could provide potential candidates for a counterexample.
@@
    In light of earlier studies on the Shannon capacity of graphs, as well as the present work, it remains unknown whether there exists a finite, undirected, and simple graph whose Shannon capacity is attained by the independence number of some finite strong power, but not by that of the first or second strong powers. Further study of this problem is therefore of interest.
  4. 4.
    By combining Theorems 2.5 and 2.6, the equality $\alpha(\mathsf{G}\boxtimes\mathsf{H})=\alpha(\mathsf{G})\,\alpha(\mathsf{H})$ holds for every simple graph $\mathsf{H}$ if $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$. Moreover, by Theorem 2.12 and Lemma 3.2, if $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$, then $\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})$ holds for all $\mathsf{H}$. From these results, a natural question arises: Is it true that the equality $\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})$ holds for all $\mathsf{H}$ if and only if $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$?. This question was already discussed to some extent in [53], showing that
@@
    while raising a question about a possible gap between the two sides of (270).
  5. 5.
    The new inequality in Theorem 7.2 can be viewed as an analogue of the arithmetic–geometric mean inequality for the Shannon capacity of graphs, relating the Shannon capacity of the strong product of graphs to that of their disjoint union. Part of the interest in this inequality arises from the fact that every finite, undirected, simple, and connected graph admits a unique prime factorization with respect to the strong product (see Theorem 7.1). Several applications of the inequality in Theorem 7.2 are presented in Section 7, and additional applications are likely to follow.
@@
Appendix A Proof of the duality result in Theorem 2.12
By Theorems 2.9 and 2.10, the claim is equivalent to
$\Theta(\mathsf{G}+\mathsf{H})>\Theta(\mathsf{G})+\Theta(\mathsf{H})\iff\Theta(\mathsf{G}\boxtimes\mathsf{H})>\Theta(\mathsf{G})\,\Theta(\mathsf{H}).$ (A.1)
We next prove both directions of the equivalence in (A.1).
  1. 1.
    Assume that $\Theta(\mathsf{G}\boxtimes\mathsf{H})>\Theta(\mathsf{G})\,\Theta(\mathsf{H})$. By equality (29),
    $\hskip-5.69046pt\Theta(\mathsf{G}+\mathsf{H})^{2}$
    $=\Theta((\mathsf{G}+\mathsf{H})^{2})$
    $=\Theta(\mathsf{G}^{2}+2\mathsf{G}\boxtimes\mathsf{H}+\mathsf{H}^{2}),$ (A.2)
    where (A.2) holds by the commutative semiring argument in [28]. By Theorems 2.9 and 2.10,
    $\Theta(\mathsf{G}^{2}+2\mathsf{G}\boxtimes\mathsf{H}+\mathsf{H}^{2})\geq\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G}\boxtimes\mathsf{H})+\Theta(\mathsf{H})^{2}.$ (A.3)
    Using the above assumption, we get
    $\hskip-5.69046pt\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G}\boxtimes\mathsf{H})+\Theta(\mathsf{H})^{2}$
    $>\Theta(\mathsf{G})^{2}+2\Theta(\mathsf{G})\,\Theta(\mathsf{H})+\Theta(\mathsf{H})^{2}$
    $=(\Theta(\mathsf{G})+\Theta(\mathsf{H}))^{2},$ (A.4)
    which gives $\Theta(\mathsf{G}+\mathsf{H})>\Theta(\mathsf{G})+\Theta(\mathsf{H})$.
  2. 2.
    $\hskip-5.69046pt\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\,\Theta(\mathsf{G})^{j}\,\Theta(\mathsf{H})^{i}$
    $=\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\,\Theta(\mathsf{G}^{j})\,\Theta(\mathsf{H}^{i})$
    $\leq\Theta(\mathsf{G}^{i+j}\boxtimes\mathsf{H}^{i+j})$
    $=\Theta(\mathsf{G}\boxtimes\mathsf{H})^{i+j}$
    $=\Theta(\mathsf{G})^{i+j}\,\Theta(\mathsf{H})^{i+j},$ (A.5)
    thus $\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\leq\Theta(\mathsf{G})^{i}\,\Theta(\mathsf{H})^{j}$, which implies that for all $k\in\mathbb{N}$
    $=\left(\Theta(\mathsf{G})+\Theta(\mathsf{H})\right)^{k}.$ (A.6)
    Finally, letting $k\to\infty$ gives $\Theta(\mathsf{G}+\mathsf{H})\leq\Theta(\mathsf{G})+\Theta(\mathsf{H})$, and by Theorem 2.10,
    $\Theta(\mathsf{G}+\mathsf{H})=\Theta(\mathsf{G})+\Theta(\mathsf{H}).$ (A.7)
Appendix B A new proof of Theorem 2.17
Theorem B.1.
[17] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then,
$\vartheta(\mathsf{G}+\mathsf{H})=\vartheta(\mathsf{G})+\vartheta(\mathsf{H}).$ (B.1)
@@
$\vartheta(\mathsf{G})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}.$ (B.2)
@@
$\vartheta(\mathsf{H})=\sum_{i=1}^{r}(\mathbf{d}^{\mathrm{T}}\mathbf{v}_{i})^{2}.$ (B.3)
Assume without loss of generality that the dimensions of $\mathbf{u}_{i},\mathbf{c},\mathbf{v}_{j}$ and $\mathbf{d}$ are identical and equal to $m$ (if the dimensions are distinct, the vectors of the lower dimension can be padded by zeros). Next, let $\mathbf{A}$ be an orthogonal matrix of order $m\times m$ such that ${\mathbf{A}}{\mathbf{d}}={\mathbf{c}}$. Such a matrix $\mathbf{A}$, satisfying ${\mathbf{A}}^{\mathrm{T}}{\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$, exists (e.g., the householder matrix defined as ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}-\frac{2(\mathbf{c}-\mathbf{d})(\mathbf{c}-\mathbf{d})^{\mathrm{T}}}{\|\mathbf{c}-\mathbf{d}\|^{2}}$ provided that $\mathbf{c}\neq\mathbf{d}$, and ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$ if $\mathbf{c}=\mathbf{d}$). Let $w=(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ be defined as $\mathbf{w}_{i}=\mathbf{A}\mathbf{v}_{i}$ for every $i\in[r]$. Since the pairwise inner products are preserved under an orthogonal (orthonormal) transformation and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ is an orthonormal representation of $\overline{\mathsf{H}}$, so is $(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$. Likewise, we get
$\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{w}_{i})^{2}=\sum_{i=1}^{r}((\mathbf{A}\mathbf{d})^{\mathrm{T}}(\mathbf{A}\mathbf{v}_{i}))^{2}=\sum_{i=1}^{r}(\mathbf{d}^{\mathrm{T}}\mathbf{v}_{i})^{2}=\vartheta(\mathsf{H}).$ (B.4)
Next, the representation $(\mathbf{x}_{1},\ldots,\mathbf{x}_{n+r})=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n},\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ is an orthonormal representation of $\overline{\mathsf{G}+\mathsf{H}}$ (since there are no additional non-adjacencies in the graph $\overline{\mathsf{G}+\mathsf{H}}$ in comparison to the disjoint union of the pairs of nonadjacent vertices in $\overline{\mathsf{G}}$ and $\overline{\mathsf{H}}$). Hence, by Theorem 2.15, the equality
$\vartheta(\mathsf{G})+\vartheta(\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{w}_{i})^{2}=\sum_{i=1}^{n+r}(\mathbf{c}^{\mathrm{T}}\mathbf{x}_{i})^{2},$ (B.5)
@@
$\vartheta(\mathsf{G})+\vartheta(\mathsf{H})\leq\vartheta(\mathsf{G}+\mathsf{H}).$ (B.6)
@@
$\vartheta(\mathsf{G}+\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{v}_{i})^{2}.$ (B.7)
@@
$\vartheta(\mathsf{G}+\mathsf{H})=\sum_{i=1}^{n}(\mathbf{c}^{\mathrm{T}}\mathbf{u}_{i})^{2}+\sum_{i=1}^{r}(\mathbf{c}^{\mathrm{T}}\mathbf{v}_{i})^{2}\leq\vartheta(\mathsf{G})+\vartheta(\mathsf{H}).$ (B.8)
Combining inequalities (B.6) and (B.8) gives the equality in (44). ∎
Appendix C The original proof of Theorem 3.4
@@
Appendix D The original proof of Theorem 5.1
Theorem D.1.
@@
$=\sum_{i=0}^{k}\sum_{j=0}^{k}\binom{k}{i}\,\binom{k}{j}\,\alpha(\mathsf{G})^{2k-i-j}\,\alpha(\mathsf{H}^{i+j}).$ (D.1)
@@
$=\sum_{i=0}^{k}\sum_{j=0}^{k}\binom{k}{i}\,\binom{k}{j}\,\alpha(\mathsf{G})^{2k-i-j}\,\alpha(\mathsf{H}^{i})\,\alpha(\mathsf{H}^{j}).$ (D.2)
Subtracting (D.2) from (D.1) gives
$\alpha(\mathsf{K}^{2k})-\alpha(\mathsf{K}^{k})^{2}=\sum_{i=0}^{k}\sum_{j=0}^{k}\binom{k}{i}\,\binom{k}{j}\,\alpha(\mathsf{G})^{2k-i-j}\,\left(\alpha(\mathsf{H}^{i+j})-\alpha(\mathsf{H}^{i})\,\alpha(\mathsf{H}^{j})\right).$ (D.3)
@@
$\alpha(\mathsf{H}^{i_{0}+j_{0}})-\alpha(\mathsf{H}^{i_{0}})\,\alpha(\mathsf{H}^{j_{0}})>0.$ (D.4)
Otherwise, by (5) and (D.4), for all $m\in\mathbb{N}$, $\alpha(\mathsf{H}^{m})=\alpha(\mathsf{H})^{m}$, so $\Theta(\mathsf{H})=\alpha(\mathsf{H})$, contradicting our assumption. Hence, $\alpha(\mathsf{K}^{2k})>\alpha(\mathsf{K}^{k})^{2}$ for all $k\geq\max\{i_{0},j_{0}\}$. Set $k_{0}\triangleq\max\{i_{0},j_{0}\}$.
Suppose by contradiction that $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ for some $k\geq k_{0}$. Then, by (5), (D.3), and (D.4), we get $\alpha(\mathsf{K}^{2k})>\alpha(\mathsf{K}^{k})^{2}$, which gives
$\Theta(\mathsf{K})\geq\sqrt[2k]{\alpha(\mathsf{K}^{2k})}>\sqrt[k]{\alpha(\mathsf{K}^{k})}=\Theta(\mathsf{K}),$ (D.5)
thus leading to a contradiction. Hence, the equality $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ cannot hold for any $k\geq k_{0}$.
Note that if we assume that the equality $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ holds for some $k<k_{0}$, then for every $n\in\mathbb{N}$
$\Theta(\mathsf{G})\geq\sqrt[nk]{\alpha(\mathsf{K}^{nk})}\geq\sqrt[k]{\alpha(\mathsf{K}^{k})}=\Theta(\mathsf{G}),$ (D.6)
which, using the same argument for $nk\geq k_{0}$, leads to a contradiction. Hence, the Shannon capacity of the graph $\mathsf{K}$ is not attained by the independence number of any finite strong power of this graph. ∎
Appendix E The original proof of inequality (245)
Theorem E.1.
@@ Appendix E The original proof of inequality (245)
$\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}.$ (E.1)
In addition, inequality (E.1) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive.
@@
  • Case 1: if there is an index $t$ that has $u_{t}\in\mathcal{A}$ and $v_{t}\in\mathcal{B}$ then obviously the two vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
  • Case 2: if there isn’t such an index, then there are two indices $1\leq i,j\leq k$, and two indices $r,s$ between $1$ and $2k$ that have $u_{r}=a_{i}$, $u_{s}=b_{i}$, and $v_{r}=a_{j}$, $v_{s}=b_{j}$. Since $\{a_{i},a_{j}\}$ is an edge in $\mathsf{G}$ if and only if $\{b_{i},b_{j}\}$ is not an edge in $\overline{\mathsf{G}}$, then the vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Hence, $\mathcal{S}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
The cardinality of $\mathcal{S}$ is $\binom{2k}{k}\,n^{k}$. Indeed, there are $\binom{2k}{k}$ ways to choose the positions of the coordinates from $\mathcal{A}$, which uniquely determines the positions of the coordinates from $\mathcal{B}$. For each such choice, there are $n$ possible values for each of the $k$ coordinates, yielding $|\mathcal{S}|=\binom{2k}{k}\,n^{k}$. Finally, we get
$\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})\geq|\mathcal{S}|=\binom{2k}{k}\,n^{k},$ (E.2)
@@
$\Theta(\mathsf{G}+\overline{\mathsf{G}})$ $\geq\lim_{k\to\infty}{\sqrt[2k]{\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})}}$
$\geq\lim_{k\to\infty}{\sqrt[2k]{\binom{2k}{k}\,n^{k}}}$
$=2\sqrt{n}.$ (E.3)
In addition, from [11, Theorem 12], inequality (E.1) indeed holds with equality when $\mathsf{G}$ is a self-complementary and vertex-transitive graph. ∎
Acknowledgments
The authors gratefully acknowledge the timely and constructive reports of the three anonymous referees, which contributed to improving the presentation.